# Quantum-secure multiparty deep learning

### v1

@@ 1 Introduction
Although deep learning has revolutionized multiple domains [1, 2, 3], its applications are constrained by increasing computational demands on the hardware [4, 5, 6]. Given the high energy consumption required for state-of-the-art deep neural network (DNN) inference [7, 8, 9], it is common to delegate inference workloads from the edge to centralized server clusters. Unfortunately, this paradigm introduces vulnerabilities that compromise data security, which is crucial in applications such as business, finance, and healthcare [4]. This situation underscores the central challenge in secure computation, where multiple parties perform a joint evaluation of multivariate functions across distributed resources while preserving the privacy of their inputs [10] (Fig. 1(a)).
Modern secure computation schemes are built on homomorphic encryption, which allows universal computation on encrypted data and preserves the security of both the inputs and the outputs of the computation [11]. Recent developments have adapted state-of-the-art homomorphic encryption schemes for secure machine learning [12], but real-world applications are limited due to their massive computational overhead [6] and recently discovered security vulnerabilities [13]. Moreover, these encryption schemes typically depend on computational complexity and are not information-theoretically secure.
In this work, we introduce a linear algebra engine for information-theoretically secure computation. We apply our engine to multiparty deep learning and derive rigorous security analysis for all parties. Applied to the MNIST classification task, we obtain test accuracies of more than $96\%$ while leaking less than $0.1$ bits per weight symbol and $0.01$ bits per data symbol. This weight leakage is an order of magnitude below the state-of-the-art minimum bit precision required for deep learning [14].
Our protocol relies on a photonic computation architecture to realize efficient optical matrix-vector multiplication [15, 16, 17]. State-of-the-art optical DNNs have been demonstrated using integrated photonics [18, 19, 20, 21, 22, 23, 24, 25, 26], free-space optics [27, 28, 29, 30, 31, 32] and fiber optics [33, 34]. Recently, a delocalized optical DNN using low-power devices was introduced and demonstrated [35, 36]. This protocol, which operates in an edge computing setting [5] where the DNN weights are streamed from a central server to clients that perform inference, shields client data from exposure but reveals the weights to the client.
To address this asymmetry, the client in our protocol performs the inference and then returns the residual light to the server as a verification state for security. Leveraging the quantum nature of light, we show that the server can compute an upper-bound on their weight leakage using the Holevo theorem [37], and that the client can upper-bound their data leakage using the Cramér–Rao bound [38].
@@ 2 Coherent linear algebra engine
      The client diverts the result mode, measures both quadratures to obtain $\vec{w}\cdot\hat{x}$. Then, the client performs a feed-forward protocol by reinjecting light with the same quadratures back into the result mode. The feed-forwarded state preserves the measured state’s expectation value $\vec{w}\cdot\hat{x}$ but, due to quantum noise, consists of additional Gaussian noise of variance $1$ SNU. The output of this measure and feed-forward step is labeled $\mathcal{M}(U_{\hat{x}}\vec{w})$.
@@ 3 Classification accuracy
In this section, we calculate the classification accuracy of a secure neural network that uses our coherent linear algebra engine on the standard MNIST classification task. To this end, we first trained a digital noiseless neural network on the MNIST dataset and got a classification accuracy of $98\%$. Then we fed the trained weights into a PyTorch model of our optical architecture to evaluate the test accuracy of our secure optical neural network, see Section 7.2 for more details.
We calculate the classification accuracy as a function of the average photon number occupation $\mu$ and the amplification gain $G$, see Section 7.2. We obtain more then $96\%$ classification accuracy with an average photon number of less then $\mu=4$ per weight and amplification gain of $G=3$.
In contrast to previous studies in optical machine learning [36, 30], we do not measure all of the light sent to the client but only the portion corresponding to $|(\vec{w}\cdot\hat{x})|^{2}$ and yet obtain similar high accuracies. This fundamentally follows from the fact that the SNR in the inner product is the similar whether one optically routes the inner product amplitude to one mode and measures it vs. measuring all the modes and calculating the inner product digitally.
@@ 4.1 Weights leakage
In this section, we use the verification state $\rho_{v}$ returned to the server by the client to calculate an upper bound on the amount of information about $\vec{w}$, in terms of number of bits, that could be learned by an arbitrary dishonest client. A dishonest client tries to learn the weights using arbitrary operations other than the operations defined by our protocol.
Here, we assume that the dishonest operations are independent and identically distributed on all the incoming modes of $\vec{w}$; the client employs separable ancilla states that interact individually with each weight mode. The ancilla states are stored in a quantum memory until the end of the attack and subsequently measured independently from one another to reveal information about $\vec{w}$. Our security analysis for the weight leakage relies on results for quantum encryption [44, 45, 46, 47, 48] and proceeds via the Holevo theorem [37]; see Section 7.3.1 for the complete analysis.
@@ 4.2 Data leakage
We recall from Section 2 that the mode $i$ of the verification state $U_{\vec{x}}^{\dagger}\mathcal{G}(U_{\vec{x}}\vec{w})$ has the mean $w_{i}$ and variance $(1+\eta_{i}$ SNU), Eq. (1). In other words, information about the client’s data $\vec{x}$ is leaked to the server not through the mean of the verification state, but through its variance. The client may control this leakage by reducing the gain $G$ in its amplification-and-splitting step. We also note that since the client is honest, it announces the true gain $G$ to the server.
We phrase the dishonest attempts of the server to learn $\vec{x}$ through $M$ measurements as the following concrete statistical estimation problem: let $\rho_{v}$, the state sent back to the server by the client, be a Gaussian random vector with mean $\vec{w}$ and a covariance matrix whose $i$-th diagonal element is $\sigma_{i}^{2}=1+\eta_{i}=1+(2-\frac{2}{G})\cdot|\hat{x}_{i}|^{2}$. What is the amount of information, in terms of bits, that the server can obtain about $\vec{x}$ from $M$ independent measurements of $\rho_{v}$, given that it knows $G$?
@@ 5 Secure classification
The optical loss present between the different parties in real-life communication networks affects the leakage calculations we presented above. We analyze the effect of loss on security in Section 7 and show that weight leakage increases with loss, while data leakage is independent of channel loss; the loss-adjusted weight and data leakage are presented in Fig. 3(a). Here we chose an average photon number of $\mu=4$ and amplification gain of $G=3$, the values we used to yield a MNIST classification accuracy of $96\%$. For local-area network or metropolitan-area networks with fiber losses of up to $6$ dB [50], the model leakage is at most $4$ bits per weight.
Since the number of neurons per layer of state-of-the-art DNNs is much larger than the size presented in our study, we present the dependence of the data and weight leakages on the number of neurons per layer in Fig.  3(b). As in pannel (a), we chose an average photon number of $\mu=4$ and amplification gain of $G=3$, the values we used to yield a MNIST classification accuracy of $96\%$. Interestingly, the leakage vanishes with the number of neurons per layer. This is a because the amplification excess noise is spread over more modes in the verification state, leading to less weight and data leakages. Therefore, advanced DNNs, which are much larger than the model used in this work, are expected to suffer from even smaller information leakage.
@@
Finally, we briefly introduce an extension of our protocol to a network consisting of a server and two clients or more, where each party assumes that all other parties are dishonest and collaborating against them. We propose two network architectures: symmetric and asymmetric.
@@ 6 Discussion
Our current protocol forces a trade-off between security for classification accuracy; future generations could relax this trade-off by joint optimization of DNN model parameters and protocol steps. Performance may be further improved by incorporating non-Gaussian operations into the protocol. For example, the client could use a photon-number-resolving detector, and the server could encode the weights using the photon number. Additionally, better performance may be achieved in a hardware-aware environment by optimizing the weights via physical custom layers instead of noiseless layers.
Our security analysis for weight and data leakage assumes individual attacks, continuous modulation of the encoding state, and asymptotically large blocks of signals per communication. Recently, however, asymptotic and finite-size security analyses of the discrete modulated continuous-variable quantum key distribution have been introduced for coherent attacks [51, 52]. These results could be applied to our quantum-secure multiparty deep learning to achieve security against other attack classes by adapting them to optimize for the security metric of leakage in our protocol.
@@
Our approach solves a fundamental security challenge in multiparty computation, paving the way for information-theoretical security at various stages of the machine learning pipeline. The introduction of our protocol enables the secure application of deep learning on private datasets and models, establishing a rigorous standard of secure computation in sectors such as finance, healthcare and business.
@@ 7.1 Optical implementation
The server consists of two modules, the transmitter and the receiver (see Fig. 4(a)). The transmitter module I/Q modulates the neural network weights onto a train of weak coherent states which are produced by attenuating a continuous wave laser to the few-photon limit. The receiver module measures the modulated quadratures of the incoming verification state using homodyne detection with a reference local oscillator. The optical power difference between the two output arms is proportional to either the I or the Q quadrature of the verification state, depending on the phase of the local oscillator.
We propose multiple optical implementations of the unitaries $U_{\hat{x}}$ and $U_{\hat{x}}^{\dagger}$ at the client. The client could operate via time domain encoding using optical loops [53] (see Fig. 4(b)). Equivalently, the client could operate by spatial domain encoding using a mesh of interferometers [54, 55] (see Fig. 4(c)) or free-space multi-plane light converters [56, 57, 58]. Time-domain implementations are promising for large-scale universal optical information processing because the number of elements does not scale with the number of optical modes [59, 60, 61]. Due to optical loss, however, both technologies are currently limited to a few dozen modes.
In the time domain design, optical modes are defined by temporally separate pulses. The first operation $U_{\hat{x}}$ is implemented by a Mach-Zehnder interferometer (MZI) and a fiber loop. This structure allows us to interfere successive pulses from the incoming pulse train $\vec{w}$ with a pulse in the loop that contains the running sum of the inner product, resulting in the final inner product $\vec{w}\cdot\hat{x}$ being written into the complex amplitude of the last output pulse. The light in this mode is then amplified by a factor of $G$ using a phase-insensitive amplifier such as a standard erbium-doped fiber amplifier. Then a beam splitter divides the light with a splitting ratio of $1-\frac{1}{G}:\frac{1}{G}$ for the transmitted and reflected ports. The transmitted port is then measured using homodyne detection in both quadratures to estimate the desired inner product. The reflected port is fed into a nested Mach-Zehnder fiber loop that implements the unitary $U^{\dagger}_{\hat{x}}$ and routes the resultant verification state back to the server.
@@ 7.1.1 Quantum description
First, we consider the effect of the overall unitary transform pair $\{U_{\hat{x}},U_{\hat{x}}^{\dagger}\}$ on the variance of the state. The initial input $\vec{w}$ is a set of coherent states with unit shot-noise variance in each mode, i.e., the noise itself is uncorrelated and isotropic. We describe the noise in these $N$ modes, and the correlations between them, via the covariance matrix of their quadratures $\Sigma^{X}$. The $(i,j)$-th element of this matrix is given by $\Sigma^{X}{(i,j)}~{}=~{}\mathbb{E}\left[X_{i}X_{j}\right]-\mathbb{E}\left[X_{i}\right]\mathbb{E}\left[X_{j}\right]$.
@@
We next apply the amplify-and-split operation $\mathcal{G}(\cdot)$ on the vector of modes $U_{\hat{x}}\vec{w}$. This operation involves applying a phase-insensitive amplification with gain $G$ on the result mode and then splitting it in the ratio $1-\frac{1}{G}:\frac{1}{G}$. The output of the $\frac{1}{G}$ port is fed back to the mesh that implements $U_{\hat{x}}^{\dagger}$. The operation $\mathcal{G}(\cdot)$ leaves all the other modes of $U_{\hat{x}}\vec{w}$ untouched. The resultant covariance matrix after amplification and splitting is the diagonal matrix $\Sigma^{X}_{\mathcal{G}(U_{\hat{x}}\vec{w})}\sim\operatorname{diag}(\sigma^{2},1,1,\ldots,1)$, where $\sigma^{2}=1+\left(2-\frac{2}{G}\right)$ is the variance of the result mode calculated in section 7.1.
@@ 7.2 Classification accuracy calculation
Separately, we trained a digital model composed of standard PyTorch layers on preprocessed MNIST images. The preprocessing of the data set included the flattening of the $28\times 28$ images to vectors of size 784, followed by centering and scaling of each vector to the range [-1, 1]. Then we trained a standard 2-layer network with 784 inputs, 784 hidden neurons, and 10 outputs, obtaining a classification accuracy of $98\%$. We then transferred these trained digital weights into the custom secure neural network and computed the accuracy achieved by the secure protocol as a function of the weight pulse energy. We found that the accuracy of the digital and analog models were in agreement up to variations caused by quantum shot noise.
@@
Note that in order to get the desired inner product, the client scales the result digitally by $\frac{\|\vec{x}\|}{\sqrt{G-1}}\frac{\|\mathbf{W}\|_{\text{RMS}}}{\sqrt{\mu}}$. However, for the purpose of deep learning, the scaling step for the client is superfluous because the client normalizes between layers regardless. We define a ”physical scaling parameter” $F$ to capture the hardware-dependent prefactor in the SNR equation,
@@
Varying $F$ changes the SNR of the inner product outputs, thereby affecting the accuracy of the secure neural network. The purpose of defining this parameter is to enable the calculation of classification accuracy as a function of the gain $G$ and average photon occupation $\mu$ without directly executing the intensive numerical calculations in the two-dimensional space spanned by these two parameters. Instead, we calculate the classification accuracy within the space spanned by $F$. The dependency of the test accuracy achieved by the secure neural network on $F$ is presented in Fig. 5. We further find that a logistic function faithfully fits the variation of the classification accuracy as a function of $F$:

### v2

@@ 1 Introduction
Although deep learning has revolutionized multiple domains [1, 2, 3], its applications are constrained by the increasing computational demands on the hardware [4, 5, 6]. Given the high energy consumption required for state-of-the-art deep neural network (DNN) inference [7, 8, 9], it is common to delegate inference workloads from the edge to centralized server clusters. Unfortunately, this paradigm introduces vulnerabilities that compromise data security, which is crucial in applications such as business, finance, and healthcare [4]. This situation underscores the central challenge in secure computation, where multiple parties perform a joint evaluation of multivariate functions across distributed resources while preserving the privacy of their inputs [10] (Fig. 1(a)).
Modern secure computation schemes are built on homomorphic encryption, which allows universal computation on encrypted data and preserves the security of the input and output of the computation [11]. Recent developments have adapted state-of-the-art homomorphic encryption schemes for secure machine learning [12], but real-world applications are limited due to their massive computational overhead [6] and recently discovered security vulnerabilities [13]. Moreover, these encryption schemes typically depend on computational complexity and are not information-theoretically secure.
In this work, we introduce a linear algebra engine for information-theoretically secure computation. We apply our engine to multiparty deep learning and derive a rigorous security analysis for all parties. Applied to the MNIST classification task, we obtain test accuracies of more than $96\%$ while leaking less than $0.1$ bits per weight symbol and $0.01$ bits per data symbol. This weight leakage is an order of magnitude below the state-of-the-art minimum bit precision required for deep learning [14].
Our protocol relies on a photonic computation architecture to realize efficient optical matrix-vector multiplication [15, 16, 17, 18]. State-of-the-art optical DNNs have been demonstrated using integrated photonics [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], free-space optics [29, 30, 31, 32, 33, 34, 35] and fiber optics [36, 37]. Recently, a delocalized optical DNN using low-power devices was introduced and demonstrated [38, 39]. This protocol, which operates in an edge computing setting [5] where the DNN weights are streamed from a central server to clients that perform inference, shields client data from exposure but reveals the weights to the client.
To address this asymmetry, the client in our protocol performs the inference and then returns the residual light to the server as a verification state for security. Leveraging the quantum nature of light, we show that the server can compute an upper-bound on their weight leakage using the Holevo theorem [40] and that the client can upper-bound their data leakage using the Cramér–Rao bound [41].
@@ 2 Coherent linear algebra engine
      The client diverts the result mode, and measures both quadratures to obtain $\vec{w}\cdot\hat{x}$. Then, the client performs a feed-forward protocol by reinjecting light with the same quadratures back into the result mode. The feed-forwarded state preserves the measured state’s expectation value $\vec{w}\cdot\hat{x}$ but, due to quantum noise, consists of additional Gaussian noise of variance $1$ SNU. The output of this measure and feed-forward step is labeled $\mathcal{M}(U_{\hat{x}}\vec{w})$.
@@ 3 Classification accuracy
In this section, we calculate the classification accuracy of a secure neural network that uses our coherent linear algebra engine on the standard MNIST classification task. To this end, we first trained a digital noiseless neural network on the MNIST dataset and obtained a classification accuracy of $98\%$. Then we fed the trained weights into a PyTorch model of our optical architecture to evaluate the test accuracy of our secure optical neural network, see Section 7.2 for more details. The complete open-source package used in this work is available online [47].
We calculate the classification accuracy as a function of the average photon number occupation $\mu$ and the amplification gain $G$, see Section 7.2. We obtain more than $96\%$ classification accuracy with an average photon number of less than $\mu=4$ per weight and amplification gain of $G=3$.
In contrast to previous studies in optical machine learning [39, 32], we do not measure all of the light sent to the client but only the portion corresponding to $|(\vec{w}\cdot\hat{x})|^{2}$ and yet obtain similar high accuracies. This fundamentally follows from the fact that the SNR in the inner product is similar whether one optically routes the inner product amplitude to one mode and measures it vs. measuring all the modes and calculating the inner product digitally.
@@ 4.1 Weights leakage
In this section, we use the verification state $\rho_{v}$ returned to the server by the client to calculate an upper bound on the amount of information about $\vec{w}$, in terms of number of bits, that could be learned by an arbitrarily dishonest client. A dishonest client tries to learn the weights using arbitrary operations other than the operations defined by our protocol.
Here, we assume that dishonest operations are independent and identically distributed over all the incoming modes of $\vec{w}$; the client employs separable ancilla states that interact individually with each weight mode. The ancillary states are stored in a quantum memory until the end of the attack and subsequently measured independently from one another to reveal information about $\vec{w}$. Our security analysis for the weight leakage relies on results for quantum encryption [48, 49, 50, 51, 52, 53] and proceeds through the Holevo theorem [40]; see Section 7.3.1 for the complete analysis.
@@ 4.2 Data leakage
We recall from Section 2 that the mode $i$ of the verification state $U_{\vec{x}}^{\dagger}\mathcal{G}(U_{\vec{x}}\vec{w})$ has the mean $w_{i}$ and variance $(1+\eta_{i}$ SNU), Eq. (1). In other words, information about the client’s data $\vec{x}$ is leaked to the server not through the mean of the verification state but through its variance. The client may control this leakage by reducing the gain $G$ in its amplification-and-splitting step. We also note that since the client is honest, it announces the true gain $G$ to the server.
We phrase dishonest attempts of the server to learn $\vec{x}$ through $M$ measurements as the following concrete statistical estimation problem[54]: Let $\rho_{v}$, the state sent back to the server by the client, be a Gaussian random vector with mean $\vec{w}$ and a covariance matrix whose $i$-th diagonal element is $\sigma_{i}^{2}=1+\eta_{i}=1+(2-\frac{2}{G})\cdot|\hat{x}_{i}|^{2}$. What is the amount of information, in terms of bits, that the server can obtain about $\vec{x}$ from $M$ independent measurements of $\rho_{v}$, given that it knows $G$?
@@ 5 Secure classification
The optical loss present between the different parties in real-life communication networks affects the leakage calculations we present above. We analyze the effect of loss on security in Section 7 and show that weight leakage increases with round-trip channel loss, while data leakage is independent of round-trip channel loss; the loss-adjusted weight and data leakage are presented in Fig. 3(a). Here we chose an average photon number of $\mu=4$ and amplification gain of $G=3$, the values we used to yield a MNIST classification accuracy of $96\%$. For local area networks or metropolitan areas networks with fiber losses of up to $6$ dB [56], the model leakage is at most $4$ bits per weight.
Since the number of neurons per layer of state-of-the-art DNNs is much larger than the size presented in our study, we present the dependence of the data and weight leakages on the number of neurons per layer in Fig.  3(b). As in pannel (a), we chose an average photon number of $\mu=4$ and an amplification gain of $G=3$, the values we used to yield a MNIST classification accuracy of $96\%$. Interestingly, the leakage vanishes with the number of neurons per layer. This is because the amplification excess noise is spread over more modes in the verification state, leading to less weight and data leakages. Therefore, advanced DNNs, which are much larger than the model used in this work, are expected to suffer even smaller information leakage.
@@
Finally, we briefly introduce an extension of our protocol to a network consisting of a server and two or more clients, where each party assumes that all other parties are dishonest and collaborate against them. We propose two network architectures: symmetric and asymmetric.
@@ 6 Discussion
Our current protocol forces a trade-off between security for classification accuracy; future generations could relax this trade-off by joint optimization of DNN model parameters and protocol steps. Additionally, better performance may be achieved in a hardware-aware environment by optimizing the weights via physical custom layers instead of noiseless layers. Performance may be further improved by incorporating non-Gaussian operations into the protocol [57, 58]. For example, the client could use a photon-number-resolving detector, and the server could encode the weights using the photon number. The use of quantum resources will also enable the extension of our protocol beyond classical computation [59], such as quantum machine learning [60, 61, 62, 63] or blind quantum computation [64, 65].
Our security analysis for weight and data leakage assumes individual attacks, continuous modulation of the encoding state, and asymptotically large blocks of signals per communication. Recently, however, asymptotic and finite-size security analyses of the discrete modulated continuous-variable quantum key distribution have been introduced for coherent attacks [66, 67, 68]. These results could be applied to our quantum-secure multiparty deep learning to achieve security against other attack classes by adapting them to optimize for the security metric of leakage in our protocol.
Recent experimental advances support the feasibility of our system. On the client side, large unitary transformations and optical computations have recently been demonstrated [69, 32, 70, 71, 72, 73]. On the server side, continuous-variable quantum key distribution has been successfully implemented [74, 75, 76]. The combination of these technologies paves the way for the near-term realization of our quantum secure multiparty deep learning.
@@
Our approach solves a fundamental security challenge in multiparty computation, paving the way for information-theoretical security at various stages of the machine-learning pipeline. The introduction of our protocol enables the secure application of deep learning on private datasets and models, establishing a rigorous standard of secure computation in sectors such as finance, healthcare, and business.
@@ 7.1 Optical implementation
The server consists of two modules, the transmitter and the receiver (see Fig. 4(a)). The transmitter module I/Q modulates the neural network weights onto a train of weak coherent states which are produced by attenuating a continuous-wave laser to the few-photon limit. The receiver module measures the modulated quadratures of the incoming verification state using homodyne detection with a reference local oscillator. The optical power difference between the two output arms is proportional to either the I or the Q quadrature of the verification state, depending on the phase of the local oscillator.
We propose multiple optical implementations of the unitaries $U_{\hat{x}}$ and $U_{\hat{x}}^{\dagger}$ in the client. The client could operate via time domain encoding using optical loops [77] (see Fig. 4(b)). Equivalently, the client could operate by spatial domain encoding using a mesh of interferometers [78, 69] (see Fig. 4(c)) or free-space multi-plane light converters [79, 80, 81]. Time-domain implementations are promising for large-scale universal optical information processing because the number of elements does not scale with the number of optical modes [82, 83, 84]. However, because of optical loss, both technologies are currently limited to a few dozen modes.
In the time domain design, optical modes are defined by temporally separate pulses. The first operation $U_{\hat{x}}$ is implemented using a Mach-Zehnder interferometer (MZI) and a fiber loop. This structure allows us to interfere successive pulses from the incoming pulse train $\vec{w}$ with a pulse in the loop that contains the running sum of the inner product, resulting in the final inner product $\vec{w}\cdot\hat{x}$ being written into the complex amplitude of the last output pulse. The light in this mode is then amplified by a factor of $G$ using a phase-insensitive amplifier such as a standard erbium-doped fiber amplifier. Then a beam splitter divides the light with a splitting ratio of $1-\frac{1}{G}:\frac{1}{G}$ for the transmitted and reflected ports. The transmitted port is then measured using homodyne detection in both quadratures to estimate the desired inner product. The reflected port is fed into a nested Mach-Zehnder fiber loop that implements the unitary $U^{\dagger}_{\hat{x}}$ and routes the resultant verification state back to the server.
@@ 7.1.1 Quantum description
First, we consider the effect of the overall unitary transform pair $\{U_{\hat{x}},U_{\hat{x}}^{\dagger}\}$ on the variance of the state. The initial input $\vec{w}$ is a set of coherent states with a unit shot-noise variance in each mode, that is, the noise itself is uncorrelated and isotropic. We describe the noise in these $N$ modes and the correlations between them, through the covariance matrix of their quadratures $\Sigma^{X}$. The $(i,j)$-th element of this matrix is given by $\Sigma^{X}{(i,j)}~{}=~{}\mathbb{E}\left[X_{i}X_{j}\right]-\mathbb{E}\left[X_{i}\right]\mathbb{E}\left[X_{j}\right]$.
@@
Next, we apply the amplify-and-split operation $\mathcal{G}(\cdot)$ on the vector of modes $U_{\hat{x}}\vec{w}$. This operation involves applying a phase-insensitive amplification with gain $G$ on the result mode and then splitting it with the ratio $1-\frac{1}{G}:\frac{1}{G}$. The output of the $\frac{1}{G}$ port is fed back to the mesh that implements $U_{\hat{x}}^{\dagger}$. The operation $\mathcal{G}(\cdot)$ leaves all the other modes of $U_{\hat{x}}\vec{w}$ untouched. The resultant covariance matrix after amplification and splitting is the diagonal matrix $\Sigma^{X}_{\mathcal{G}(U_{\hat{x}}\vec{w})}\sim\operatorname{diag}(\sigma^{2},1,1,\ldots,1)$, where $\sigma^{2}=1+\left(2-\frac{2}{G}\right)$ is the variance of the result mode calculated in Section 7.1.
@@ 7.2 Classification accuracy calculation
Separately, we trained a digital model composed of standard PyTorch layers on preprocessed MNIST images. The preprocessing of the data set included the flattening of the $28\times 28$ images to vectors of size 784, followed by a centering and scaling of each vector to the range [-1, 1]. Then we trained a standard 2-layer network with 784 inputs, 784 hidden neurons, and 10 outputs, obtaining a classification accuracy of $98\%$. We then transferred these trained digital weights into the custom secure neural network and computed the accuracy achieved by the secure protocol as a function of the weight pulse energy. We found that the accuracy of the digital and analog models were in agreement up to variations caused by quantum shot noise.
@@
Note that in order to obtain the desired inner product, the client scales the result digitally by $\frac{\|\vec{x}\|}{\sqrt{G-1}}\frac{\|\mathbf{W}\|_{\text{RMS}}}{\sqrt{\mu}}$. However, for the purpose of deep learning, the scaling step for the client is superfluous because the client normalizes between layers regardless. We define a ”physical scaling parameter” $F$ to capture the hardware-dependent prefactor in the SNR equation,
@@
Varying $F$ changes the SNR of the inner product outputs, thus affecting the accuracy of the secure neural network. The purpose of defining this parameter is to enable the calculation of the classification accuracy as a function of the gain $G$ and average photon occupation $\mu$ without directly executing the intensive numerical calculations in the two-dimensional space spanned by these two parameters. Instead, we calculate the classification accuracy within the space spanned by $F$. The dependency of the test accuracy achieved by the secure neural network on $F$ is presented in Fig. 5. We further find that a logistic function faithfully fits the variation of the classification accuracy as a function of $F$: